# Impact of general relativistic accretion on primordial black holes

### v1

@@ VI Improved constraints on DM
An important constraint on the DM parameter space comes from the warm dark matter (WDM) bound. PBHs can emit DM particles with high momenta, which may remain relativistic until structure formation, conflicting with observations. This bound is given by Haque:2023awl ; Barman:2024iht ,
@@
In this analysis, we adopt the lower bound $m_{\rm WDM}>3.3\,\rm keV$ Viel:2013fqw ; Bode:2000gq . This constraint is shown in yellow in Fig. 3.
@@ VIII Gravitational Waves from PBH Evaporation
Relativistic accretion extends PBH lifetimes by modifying the mass loss rate, which in turn shifts the peak frequency and amplitude of the resulting GW spectrum (see Fig.  5). Detecting these high-frequency GWs is challenging but potentially feasible with next-generation detectors, including laser interferometers (1–10 kHz)  Ackley:2020atn , levitated sensors (10–100 kHz)  Aggarwal:2020olq , magnetic conversion (10 GHz)  Li:2009zzy ; Ito:2019wcb , and inverse Gertsenshtein effect devices ($10^{14}–10^{18}$ Hz)  Ejlli:2019bqj . Axion experiments such as JURA Beacham:2019nyx , OSQAR II OSQAR:2015qdv , and CAST  CAST:2017uph may be repurposed for graviton detection, providing novel probes of PBH accretion and early-universe dynamics.
@@ IX Conclusions
PBHs have become an integral part of the theoretical model building of the early Universe and dark matter. In this Letter, we revisit the cosmological evolution of PBHs by incorporating two key physical processes, namely Hawking evaporation and accretion. In contrast to previous studies that modeled accretion using a pseudo-Newtonian approximation, typically yielding negligible mass growth, we formulate the accretion dynamics within a fully general relativistic framework. We find that relativistic corrections significantly enhance the accretion rate, leading to rapid PBH mass growth on timescales much shorter than those of Hawking evaporation. This enhancement imposes a new lower bound on the PBH mass, $M_{\rm in}\leq 6.9\times 10^{7}$ g, above which excessive energy injection from evaporation can disrupt BBN.
Furthermore, we analyze the consequences of this enhanced accretion on DM production via Hawking radiation and find a noticeable shift in the allowed parameter space spanned by PBH abundance and DM mass, modifying constraints relative to previous results Cheek:2021odj ; Gondolo:2020uqv . We also determine that the minimum PBH mass capable of surviving to the present day is reduced to $\sim 2\times 10^{14}$ g, which is an order of magnitude below existing estimates Carr:2009jm ; Carr:2020xqk . PBHs with masses above this threshold could themselves constitute viable DM candidates. Notably, a broad mass window between $10^{16}\,{\rm g}$ and $10^{21}\,{\rm g}$ remains observationally unconstrained, allowing PBHs within this range to account for the entirety of DM. Finally, we compute modifications to the high-frequency GW spectrum arising from accretion-driven PBH mass growth, offering a potential observational signature for future detectors.

### v2

@@ VI Improved constraints on DM
An important constraint on the DM parameter space comes from the warm dark matter (WDM) bound. PBHs can emit DM particles with high momenta, which may remain relativistic until structure formation, conflicting with observations. The present day average DM velocity, expressed in terms of the average DM momentum $\langle p_{x}\rangle$, is given by
$v_{0}=\frac{a_{\rm ev}}{a_{0}}\frac{\langle p_{x}\rangle}{m_{x}}=\frac{a_{\rm ev}}{a_{i}}\frac{1}{1+z_{\rm eq}}\frac{T_{\rm eq}}{T_{\rm acc}}\frac{\langle p_{x}\rangle}{m_{x}},$ (32)
where $a_{\rm ev}$ and $a_{0}=1$ denote the scale factors at PBH evaporation and the present epoch, respectively. Here, $z_{\rm eq}\approx 3400$ is the redshift at matter-radiation equality, and the corresponding radiation temperature is $T_{\rm eq}\simeq 0.8\times 10^{-9}$ GeV. For light DM particles $m_{x}\ll T_{\rm BH}^{\rm in}$, the average momentum is approximately $\langle p_{x}\rangle\sim T_{\rm BH}^{\rm in}$. The present-day velocity of WDM is given by Bode:2000gq
$v_{\rm WDM}\simeq 3.9\times 10^{-8}\left(\frac{\rm KeV}{m_{\rm WDM}}\right)^{\frac{4}{3}}.$ (33)
In our present analysis, we impose the lower bound on the WDM mass to be, $m_{\rm WDM}>3.3\,\rm KeV$ Viel:2013fqw . This, in turn, constrains the DM mass as Haque:2023awl ; Barman:2024iht
@@
This constraint is shown in yellow in Fig. 4. Another important constraint arises from inflation. As discussed earlier, the Hubble parameter at the time of PBH formation is related to its initial mass by $H_{\rm in}=4\pi\gamma M_{p}^{2}M_{\rm in}^{-1}$. At the end of inflation, the Hubble parameter typically reaches a minimum value of $H_{\rm end}\sim 10^{14}\,\mathrm{GeV}$ ParticleDataGroup:2024cfk , which imposes a lower bound on the PBH mass of approximately $M_{\rm in}\gtrsim 0.1\,\mathrm{g}$. This constraint is represented by the light blue shaded region in Fig. 4.
To account for the effects of evaporation products on BBN within the mass range $10^{8}\,\mathrm{g}\lesssim M_{\rm BH}\lesssim 10^{13}\,\mathrm{g}$, we follow the framework of Ref. Keith:2020jww , which models the photodisintegration rates of light nuclei. This constraint is represented by the green shaded region in Fig. 4. We further incorporate our modified evaporation history, which includes the impact of enhanced accretion, shown by the light green shaded region in the same figure.
@@ VIII Gravitational Waves from PBH Evaporation
Relativistic accretion extends PBH lifetimes by modifying the mass loss rate, which in turn shifts the peak frequency and amplitude of the resulting GW spectrum (see Fig.  6). While our primary results and the analytical expressions derived here rely on the assumption of a monochromatic PBH mass function, we note that an extended mass function, $\psi(M_{\rm in})$, would modify the resulting spectrum. In this generalized case, the total observable GW spectrum is the superposition of all the monochromatic spectra, weighted by the mass distribution $\psi(M_{\rm in})$. Since the peak frequency scales directly with mass Ireland:2023avg , the primary effect would be a broadening of the spectrum. Consequently, the single, sharp peak corresponding to a single $M_{\rm in}$ would be “smeared out” over a range of frequencies defined by the width and shape of the PBH mass function itself.
Detecting these high-frequency GWs is challenging but potentially feasible with next-generation detectors, including laser interferometers (1–10 kHz)  Ackley:2020atn , levitated sensors (10–100 kHz)  Aggarwal:2020olq , magnetic conversion (10 GHz)  Li:2009zzy ; Ito:2019wcb , and inverse Gertsenshtein effect devices ($10^{14}–10^{18}$ Hz)  Ejlli:2019bqj . Axion experiments such as JURA Beacham:2019nyx , OSQAR II OSQAR:2015qdv , and CAST  CAST:2017uph may be repurposed for graviton detection, providing novel probes of PBH accretion and early-universe dynamics.
@@ IX Conclusions
Primordial Black Holes (PBHs) have become an integral part of theoretical model building for the early Universe and dark matter. In this paper, we revisit the cosmological evolution of PBHs by incorporating two key physical processes: Hawking evaporation and accretion. In contrast to previous studies that modeled accretion using a pseudo-Newtonian approximation, which typically yields negligible mass growth, we formulate the accretion dynamics within a fully general relativistic framework. We demonstrate that relativistic corrections, crucial in the dense plasma of the early universe, significantly enhance the accretion rate, leading to a phase of rapid, runaway PBH mass growth on timescales much shorter than those of Hawking evaporation. This dynamic interplay imposes a new and stronger lower bound on the PBH mass, $M_{\rm in}\leq 6.9\times 10^{7}\,\rm g$, as PBHs forming below this mass would grow so efficiently that their subsequent, more powerful evaporation injects excessive energy into the primordial plasma, disrupting the successful predictions of Big Bang Nucleosynthesis (BBN).
This accretion-dominated evolution has several profound cosmological implications that we have explored. We analyze the consequences of this enhanced accretion on Dark Matter (DM) production via Hawking radiation and find a noticeable shift in the allowed parameter space spanned by the PBH abundance and DM mass, modifying constraints relative to previous results Cheek:2021odj ; Gondolo:2020uqv . This modification arises because the rapid mass increase alters the evaporation lifetime and temperature of the PBHs ($T_{\rm BH}\propto 1/M_{\rm BH}$), thereby changing the production efficiency and viable parameter space for DM particles generated via Hawking radiation. As a PBH grows, it ceases to produce massive DM particles sooner than a non-accreting PBH would, thus altering the final relic abundance.
Furthermore, we determine that the minimum PBH mass capable of surviving to the present day is reduced to $\sim 2\times 10^{14}\,\rm g$, an order of magnitude below existing estimates Carr:2009jm ; Carr:2020xqk . PBHs with masses above this threshold could themselves constitute viable DM candidates. This result significantly revises the minimum mass for which a PBH can be considered a potential dark matter constituent today, effectively widening the available parameter space. Notably, our work reinforces the viability of the broad mass window between $10^{16}\,\rm g$ and $10^{21}\,\rm g$, which remains observationally unconstrained and allows PBHs to account for the entirety of DM. Finally, we compute modifications to the high-frequency gravitational wave (GW) spectrum arising from this accretion-driven PBH mass growth.
In summary, our work underscores the critical importance of incorporating a fully general relativistic treatment for PBH accretion. By doing so, we have refined key cosmological constraints related to BBN and dark matter, and we have highlighted a clear, potentially observable signature in the GW background. Future multi-messenger explorations of the early Universe will be crucial for testing the accretion-driven scenario presented herein. Important directions in which our present work can be extended could to take into account Kerr-black holes, impact of thermal feedback on the accretion process etc.